OpenStudy (anonymous):

Choose the correct simplification of the expression -6x^2(4x - 7x^2 - 2). 42x^4 + 24x^3 - 12x^2 -42x^4 + 24x^3 - 12x^2 -13x^4 - 2x^3 - 8x^2 42x^4 - 24x^3 + 12x^2

OpenStudy (cwrw238):

distribute -6x^2 over each term in the parentheses that is multiply each term in the parentheses by -2x^2

jagr2713 (jagr2713):

Its D

OpenStudy (cwrw238):

remember when you are multiplying the terms in x you ADD the powers eg x^2 * x = x^(2+1) = x^3 [note x = x^1]

OpenStudy (cwrw238):

@jagr2713 - please just don't give the answer - read Code of Conduct

OpenStudy (anonymous):

I'm sorry, I'm so confused lol.I'm following along but there's no way I could do this by myself.

OpenStudy (cwrw238):

i'll start it off for you -6x^2(4x - 7x^2 - 2) = -6^2 * 4x - 6x^2* -7x^2 - 6x^2 * -2 the first multiplication is -24 x^(2+1) = -24x^3 are you ok with that? remember that a negative * positive = negative:- 6*-4 = -24

OpenStudy (cwrw238):

* error in third line- it should be -6x^2 not -6^2

OpenStudy (anonymous):

Wait, so did you just distribute ?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

yes and worked the first part out for you

OpenStudy (anonymous):

Okay, and then next step was multiplying the terms. Now what's the next part ??

OpenStudy (cwrw238):

-6x^2 * -7x^2 = -7*-6 x^(2+2) = 42 x^4

OpenStudy (cwrw238):

now can you work out -6x^2 * -2 ?

OpenStudy (cwrw238):

after multiplying out the 3 parts just put them together

OpenStudy (anonymous):

I feel so stupid D; But no, can you help me ?

OpenStudy (anonymous):

How do we put them together ?..

OpenStudy (cwrw238):

well you just multiply the numbers first 6 * -2 = ?

OpenStudy (anonymous):

-12 ?

OpenStudy (cwrw238):

yes thats right the last term = -12x^2 so putting them together you get -24x^3 + 42x^4 - 12x^2

OpenStudy (cwrw238):

now you can rearrange these as long as you take the sign with you this is the same as 42x^4 - 24x^3 - 12x^2 which is the last option

OpenStudy (cwrw238):

jagr had the right answer but didnt explain how to get it

OpenStudy (anonymous):

Thank you so much for working through this with me.

OpenStudy (cwrw238):

yw hope you have learned something

OpenStudy (cwrw238):

OH hold on - theres an error the last term is -6x^2 * -2 = + 12x^2

OpenStudy (cwrw238):

thats still D

OpenStudy (cwrw238):

-6 * -2 = +12

OpenStudy (cwrw238):

somewhere along the line -6 became 6 - my error